Hrubec Products, Inc., operates a Pulp Division that manufactures wood pulp for use in the production of various paper goods.Complete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3 Req 4A Reg 4B Reg 5 Reg 6 What is the loComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3 Req 4A Reg 4B Reg 5 Reg 6 If the Pulp DiComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3 Req 4A Req 4B Reg 5 Reg 6 What is the lo1. What is the lowest acceptable transfer price from the perspective of the Pulp Division? What is the highest acceptable traComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3 Reg 4A Reg 4B Req 5 Reg 6 Refer to (4) a
Answer #1

Lowest Acceptable transfer price for Pulp Division = $21 i.e.Selling Price to outside customers

Highest Acceptable transfer price for Carton Division = $21*0.9 = $18.9 (current purchase price)

No Acceptable Range

Not likely to Agree

2.Effect on pulp Division:

Loss on 31,000 tons = (21-18.9)*31,000 = $65,100

Profits of pulp division will decrease by $65,100

Carton Division = No effect (same price as before)

Company as a whole = Loss of $65,100

i.e. profits will decrease by $65,100

3.Space Capacity Exists

Lowest Acceptable transfer price = Variable Cost = $11

Highest Acceptable Transfer Price for Carton Division = $21*0.9 = $18.9 (current purchase price)

Range of acceptable transfer price = $11-$18.9

Yes, they are likely to voluntarily agree

4a – The price is still higher than the variable cost to the pulp division, it should meet this price.

4b Effect on profits of the company is equal to loss of $5 per ton

=31,000*$5 = loss of $155,000

i.e. profits will reduce by $155,000

5. The carton division can accept a higher price for the good of the company as a whole

6. Effect on profits of the company:

Increase in profits of pulp division = 31,000*(21-11) = $310,000

Loss of Carton Division = 31,000*(21-18.9) = $65,100

Net benefit to the company = $244,900

Hence, profits will increase by $244,900